# CL-RAG: Bridging the Gap in Retrieval-Augmented Generation with Curriculum Learning

### v1

@@ Appendix C Complete Results of Robustness Test
EM F1 EM F1 EM F1 EM F1 EM F1
$With$ $irrelevant$ $documents$
$Llama3_{8B}$ 38.56 51.72 71.28 80.61 42.68 48.85 23.89 37.99 44.10 54.79
$RAAT$ 30.31 38.65 69.34 74.61 35.90 39.29 21.13 28.96 39.17 45.38
$RALM_{golden}$ 41.49 50.92 78.52 83.10 46.97 49.44 30.46 40.80 49.36 56.07
$RALM_{top5}$ 46.32 57.35 78.66 83.52 53.60 56.33 35.32 47.01 53.48 61.05
$RetRobust$ 46.56 57.17 79.57 84.42 54.28 57.17 35.00 46.95 53.85 61.43
$RALM_{CL}$ 50.30 59.89 81.48 86.31 56.13 59.05 36.75 48.43 56.17 63.42
$With$ $counterfactual$ $documents$
$Llama3_{8B}$ 38.10 50.78 72.53 80.93 42.49 48.52 24.16 37.95 44.32 54.55
$RAAT$ 30.60 39.16 70.23 75.40 34.92 38.59 21.32 29.43 39.27 45.65
$RALM_{golden}$ 40.08 49.03 77.52 81.90 45.58 48.07 29.82 39.92 48.25 54.73
$RALM_{top5}$ 46.18 56.88 79.10 83.88 53.88 56.66 35.34 47.10 53.63 61.13
$RetRobust$ 46.56 57.47 80.39 85.13 53.90 57.04 35.80 47.42 54.16 61.77
$RALM_{CL}$ 50.16 59.82 81.74 86.55 55.89 58.68 37.04 48.82 56.21 63.47
Table 6: Complete results on robustness test(%).
Appendix D Further Study of Retriever Training
D.1 Complete Results of Retriever Training
Complete document retrieval results are shown in Table 7 and Table 8. The final results on QA datasets are provided in Table 9.
Method NQ TriviaQA
R@1 R@5 R@10 R@1 R@5 R@10
$Baseline$ 43.64 72.75 81.00 65.56 88.25 93.30
$Replug$ 48.43 75.54 83.22 66.50 88.59 93.13
$w/o\ RALM$ 45.29 73.60 81.32 56.13 80.26 86.29
$LLM$-$Embedder$ 46.60 73.56 82.16 68.20 88.68 92.47
$w/o\ RALM$ 44.77 72.01 80.12 67.41 88.27 92.90
$CL$$-RAG$ 48.82 75.33 83.33 68.59 89.86 93.91
$w/o\ RALM$ 46.11 72.26 80.61 69.26 89.00 93.22
Table 7: Retrieve results of Natural Questions and TriviaQA.
Method PopQA HotpotQA
R@1 R@5 R@10 R@1 R@5 R@10
$Baseline$ 43.03 66.51 73.56 25.16 42.07 49.14
$Replug$ 54.79 80.50 86.15 26.68 44.63 51.23
$w/o\ RALM$ 50.55 76.49 83.90 23.30 38.64 45.48
$LLM$-$Embedder$ 54.68 79.17 85.66 28.30 44.52 50.82
$w/o\ RALM$ 51.54 75.72 82.05 27.41 43.13 50.07
$CL$$-RAG$ 55.26 81.53 87.18 28.91 46.45 53.79
$w/o\ RALM$ 51.84 78.04 84.77 26.98 43.55 50.73
Table 8: Retrieve results of PopQA and HotpotQA.
Method NQ TriviaQA PopQA HotpotQA Avg.
EM F1 EM F1 EM F1 EM F1 EM F1
$LLM$-$Emebedder$ 52.10 61.53 83.03 87.72 65.24 67.88 38.86 50.91 59.81 67.01
$Replug$ 52.73 62.42 83.17 87.67 66.16 68.75 38.73 50.33 60.20 67.29
$CL$-$RAG$ 53.01 62.51 82.97 87.57 66.63 69.12 39.11 51.19 60.43 67.60
Table 9: Final results on QA datasets when all retrievers are trained to align with the preferences of $RALM_{CL}$.(%)
D.2 Case Study of Documents Evaluation
@@
Table 10 presents the case study comparing the document evaluation result of the base LLM and the RALM. The result indicate that for some documents containing implicit knowledge required for answering questions, the base LLM may generate erroneous judgments, whereas the trained RALM is more capable of distinguishing the quality of documents. This also highlights the necessity of the iterative process of training the RALM first and then training the retriever, as the trained RALM and the base LLM may no longer share the same preferences for documents. Separate training may lead to suboptimal final outcomes.
Question: Who scored a film based on a 1961 science fiction novel by Stanislaw Lem?
@@
(Poland, Germany, and the Soviet Union). Franz Rottensteiner, Lem’s former agent abroad, had
this to say about Lem’s reception on international markets: His best-known novels include "Solaris"
(1961), "His Master’s Voice" ("Głos pana", 1968), and the late "Fiasco" ("Fiasko", 1987). "Solaris"
was made into a film in 1968 by Russian director Boris Nirenburg, a film in 1972 by Russian director
Andrei Tarkovsky—which won a Special Jury Prize at the Cannes Film Festival in 1972—and
an American film in 2002 by Steven Soderbergh. "Solaris" is not the only work of Lem’s
@@
Table 10: An example of document evaluation.
Appendix E Complete Results of Ablation Study
Complete results of the ablation study are shown in Table 11.
Method NQ TriviaQA PopQA HotpotQA Avg.
EM F1 EM F1 EM F1 EM F1 EM F1
$With$ $RALM$ $Training$
$w/o\ CL$ 49.31 59.37 80.34 85.38 57.65 60.47 35.11 46.94 55.60 63.04
$stage_{1}$ 46.32 56.26 81.77 85.99 53.29 55.79 32.77 43.79 53.54 60.46
$stage_{2}$ 51.22 60.73 83.12 87.50 58.54 61.52 37.36 49.13 57.56 64.72
$RALM_{CL}$ 51.53 61.19 82.93 87.59 59.71 62.28 38.14 50.22 58.08 65.32
$With$ $Retriever$ $Training$
$w/o\ CL$ 52.13 61.42 82.48 87.11 64.26 66.70 38.20 50.13 59.27 66.34
$stage_{1}$ 52.66 62.11 82.96 87.67 63.67 66.14 38.94 50.89 59.56 66.70
$stage_{2}$ 52.98 62.39 83.04 87.49 65.10 67.53 39.05 51.26 60.04 67.17
$CL$-$RAG$ 53.01 62.51 82.97 87.57 66.63 69.12 39.11 51.19 60.43 67.60
Table 11: Complete results of ablation study(%).

### v2

@@ IV-C Baselines
[…]
$Llama3_{8B}$
naivenaive-RAGRAG 39.51 53.37 72.74 81.80 46.07 52.16 25.23 39.46 45.89 56.70
WithWith RALMRALM TrainingTraining
RAATRAAT[14] 31.65 40.39 71.84 77.02 38.49 41.87 22.16 30.26 41.04 47.39
$RALM_{golden}$[13] 42.90 52.30 79.55 83.84 50.06 52.45 31.25 41.96 50.94 57.64
$RALM_{top5}$[7] 48.15 58.95 80.32 85.05 57.58 60.13 36.57 48.46 55.66 63.15
RetRobustRetRobust[12] 47.83 58.60 81.25 85.96 57.51 60.31 37.02 48.99 55.90 63.47
WithWith RetrieverRetriever TrainingTraining
RARA-DITDIT[13] 49.52 59.94 76.62 81.10 61.73 64.47 36.14 47.82 56.00 63.33
ReplugReplug[11] 52.52 62.28 76.86 81.59 63.75 66.35 37.75 49.22 57.72 64.86
LLMLLM-EmbedderEmbedder[21] 51.85 61.41 82.60 87.20 64.24 66.69 38.80 50.51 59.37 66.45
DACLDACL-$RAG(Ours)$ 53.01 62.51 82.97 87.57 66.63 69.12 39.11 51.19 60.43 67.60
$Qwen2.5_{7B}$
naivenaive-RAGRAG 37.65 50.93 74.73 80.65 46.22 51.21 29.38 39.69 47.00 55.62
WithWith RALMRALM TrainingTraining
RAATRAAT[14] 30.73 39.67 74.24 79.85 37.96 42.59 27.94 35.47 42.72 49.40
$RALM_{golden}$[13] 41.06 50.20 75.93 80.78 48.30 51.04 30.71 41.34 49.00 55.84
$RALM_{top5}$[7] 44.80 55.06 75.28 81.04 55.38 58.33 35.21 46.41 52.67 60.21
RetRobustRetRobust[12] 45.05 55.35 77.09 82.78 54.93 57.66 35.77 47.45 53.21 60.81
$RALM_{CL}(Ours)$ 48.47 58.29 79.44 84.67 55.89 59.10 37.38 49.05 55.30 62.78
WithWith RetrieverRetriever TrainingTraining
RARA-DITDIT[13] 44.06 53.54 75.89 80.77 58.45 61.18 31.45 42.07 52.46 59.39
ReplugReplug[11] 49.67 59.90 78.95 84.34 62.53 65.68 35.32 47.10 56.62 64.26
LLMLLM-EmbedderEmbedder[21] 49.45 59.22 79.08 84.34 60.53 63.35 35.86 47.61 56.23 63.63
DACLDACL-$RAG(Ours)$ 51.11 60.56 80.34 85.48 63.56 66.57 36.48 48.47 57.87 65.27
IV-D Implementation Details
IV-D1 RALM Training
We trained our model based on LLaMA3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct[3] and Qwen2.5-7B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct[54], using five retrieved documents. We randomly selected 6000 training samples, 1500 from each of four datasets, and conducted LoRA finetuning based on LLaMA Factory [^4] https://github.com/hiyouga/LLaMA-Factory[55], with the LoRA rank set to 8, a learning rate of 1e-4, a gradient accumulation step of 8, and a warmup ratio of 0.1. All experiments were conducted on an A800 80G GPU card.
IV-D2 Retriever Training
We trained our model based on Contriever-MSMARCO [52], a widely used unsupervised dense retriever with strong cross-dataset transfer and broad domain generality. By randomly sampling from the training sets of four datasets, we constructed a retriever training set comprising 100,000 samples. We set the number of input retrieved documents $k$ to 5, the learning rate of 1e-5, the number of epochs to 3, and the batch size to 64. For our three-stage curriculum learning settings, the number of reranked documents $n$ is set to 20. $n_{1}=[1,3,5]$ for each training epoch and $n_{2}$ is set consistently to 15. The sampling parameters for each stage are $k_{1}=[1,3,5]$, $k_{2}=[2,2,0]$, and $k_{3}=[2,0,0]$.
IV-D3 Inference
During inference, we used the same prompt across all models to ensure fairness, as shown in Table IV. In this prompt, paragraph represents the retrieved document, and instruction represents the query. We extracted the content following ”Answer:” in the model’s response as the final answer.
V Results
V-A Main Results
The overall results are shown in Table V, where we compare our method with baselines across two distinct model families.
V-A1 With RALM Training
Results in the upper half of Table V show that our method outperforms prior approaches across all test sets on both model backbones, yielding improvements of 2.18% in average EM and 1.85% in average F1 over the best baseline, thereby validating the effectiveness and generalization of our approach. It is noted that the model trained with the RAAT method performs worse than the baseline model. This is because its training data follows the setting of $\{d_{golden}\oplus d_{noise}\}$. While this setting trains the model to be robust against noise, the inclusion of golden documents in all data deviates from the conventional RAG setup, resulting in poorer performance under the standard setting. Training using documents that contain a golden document can enhance the model’s basic performance, i.e., its ability to extract correct answers from documents. However, due to the limitations of the retriever, the retrieved documents may not always contain the necessary information to answer the question. Therefore, training solely for answer extraction is insufficient. It is also necessary to further train the model to infer answers from more complex documents. RetRobust considers a more complex setting by accounting for both irrelevant document noise and counterfactual document noise. However, it neglects the incorporation of golden documents, which are beneficial in the early stages of model training. In contrast, our proposed method considers all types of document combinations, thereby enabling the model to adapt to the diverse documents generated by retrieval. Therefore, it achieves the best performance.
V-A2 With Retriever Training
In the settings that incorporate retriever training, our method also achieves the best average performance. As shown in the lower half of Table V, RA-DIT utilizes the model from its experimental setting, namely $RALM_{top5}$, while the other methods all employ the best-performing $RALM_{CL}$. We found that on datasets where LLM itself performs well, such as TriviaQA, the gains from further improving the retriever’s performance are minimal. However, for datasets like PopQA, which involve questions about specific entities, enhancing the retriever enables it to more accurately fetch entity-related documents, resulting in a significant boost in performance.
V-A3 Robustness Test
TABLE VI: Experimental results on robustness test(%) over irrelevant and counterfactual documents. We report the average results on each dataset.
Method Irrelevant Counterfactual
$uparrow EM\uparrow F1\uparrow$
$Llama3_{8B}$ 44.10 54.79 44.32 54.55
RAATRAAT 39.17 45.38 39.27 45.65
$RALM_{golden}$ 49.36 56.07 48.25 54.73
$RALM_{top5}$ 53.48 61.05 53.63 61.13
RetRobustRetRobust 53.85 61.43 54.16 61.77
$RALM_{CL}$ 56.17 63.42 56.21 63.47
To demonstrate the superiority of our RALM in combating noise, we also conducted tests under document settings that included irrelevant noise or counterfactual noise. The results are shown in Table VI. Under the influence of document noise, the performance of all models declined to some extent. The impact of irrelevant document noise was comparable to that of counterfactual noise. The experimental results show that our model maintained its leading performance even when different types of noise were added.
V-B Generalization Studies
Figure 4: Generalization experiments of DACL-RAG training methods across Qwen2.5-3B model, we report the average Exact Match (EM) scores for generation accuracy evaluation and Recall@k scores for retrieval quality assessment, which demonstrate consistent superiority of our approach.
In this section, we will examine the generalization capability of our proposed DACL-RAG training framework, focusing on two key aspects: (1) the generalization of the method when applied to different models, and (2) the generalization of models trained with our method on out-of-domain datasets.
V-B1 Generalization experiment of different models
To evaluate the generalizability of our method across different model parameter scales, we conducted additional experiments using Qwen2.5-3B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-3B-Instruct models[54]. We tested the performance of generator training as well as the retrieval training guided by this model. The key findings are presented in Figure 4.
For generator training, We compare our method against the strongest baseline (RetRobust), with all methods using identical documents retrieved by base retriever Contriever-MSMARCO. The results show that our method still maintains a performance advantage across different model parameters, achieving an average improvement of 10.15% in EM scores compared to the baseline, and an average performance gain of 2.15% over the strong baseline RetRobust. This indicates that our method framework generalizes across different model parameters and can be applied to models with varying parameter sizes while maintaining consistent leading performance.
For retriever training, we employ the corresponding well-trained models to score and rank documents, comparing our method against the strong baseline (LLM-Embedder). All generators use corresponding models trained with our DACL-RAG method. Our approach demonstrates superior performance in both document recall rate and final answer accuracy compared to the baselines. These results indicate consistent improvements across different model sizes, confirming the robust generalization capability of our method.
V-B2 Generalization experiment of out-of-domain dataset
To test the generalization ability of our method on unseen data, we conducted evaluations using an out-of-domain dataset. Specifically, we further assessed the performance of our trained retriever and generator on the 2Wiki[51] dataset (excluded from training), comparing it with strong baseline methods LLM-Embedder and RetRobust. The results are shown in Figure 5, where EM and F1 scores reflect the performance of the final generated answers, while R@k measures the retriever’s retrieval performance.
The experimental results demonstrate that our method achieves strong performance even on untrained datasets, outperforming the baseline approaches. We observed that in terms of document retrieval recall, while LLM-Embedder shows a slight improvement in R@1, its R@5 performance actually decreases by 1% compared to the original retriever, which also negatively impacts its final answer generation performance. In contrast, our method demonstrates excellent performance in overall recall rates, improving R@1 by 1.54% while simultaneously increasing R@5 by 2.68%, delivering consistent performance improvements. This indicates that our method achieves better results than previous methods in retrieving multiple relevant documents, while also demonstrating the excellent generalization capability of our method.
V-C Ablation Studies
In this section, we will discuss each component of our DACL-RAG framework, including: (1) research on document rewriting, (2) in-depth study on retriever training, and (3) ablation study of each training stage. In these experiments, we select the aforementioned four main datasets: NQ, TriviaQA, PopQA, and HotpotQA, using the LLaMA3-8B-Instruct model as the base generator and Contriever-MSMARCO as the base retriever.
Figure 5: Evaluation results on the out-of-domain dataset 2WikiMultiHopQA, where EM and F1 assess answer accuracy, and R@k evaluates the quality of retrieved documents.
V-C1 Research on Document Rewriting
Figure 6: Performance comparison between training with golden documents only versus combined golden and rewritten documents. The joint training approach demonstrates consistent advantages across all test sets.
Figure 7: Comparative analysis of data quality across different rewrite models at different training stages, we report average results for each training dataset, revealing quality differences between Llama and Qwen models in document rewriting.
For the document rewriting strategy proposed in our method, we conducted further research in two aspects: whether using both rewritten documents and golden documents for training actually enhances model performance compared to using only golden documents, and a study on rewriting quality when using different models.
For the first question, we compared the performance differences when training with only 5 documents containing golden documents versus training with 5 documents containing both rewritten golden documents and original golden documents. The results are shown in Figure 6. The results show that compared to the method using only golden documents for training, our method has advantages across all datasets, bringing an average improvement of 2.67% in EM score. This indicates that in data already containing golden documents, adding rewritten documents that also contain correct answers, thereby ensuring the document set contains two or more documents with correct answers, can effectively enable the model to learn the ability to select documents with higher consistency from multiple different documents for answering, thus improving performance.
Regarding rewriting document quality, we compared the rewriting quality based on Llama model and Qwen model (here we used the Qwen2.5-7B-Instruct model), and evaluate the impact of different training stages on LLM generation accuracy and document recall rate. The results are shown in Figure 7, where for common difficulty the documents are not rewritten and thus share the same recall rate, while for easy difficulty the top five documents are guaranteed to contain the correct answer, resulting in 100% R@5. Our analysis reveals that Llama demonstrates superior rewriting performance compared to Qwen across both document types: it achieves higher average recall on first-stage training data and lower average recall on third-stage training data, indicating better performance in both golden document rewriting and fake document rewriting. This suggests Llama may be the more optimal choice for rewriting tasks.
V-C2 In-depth Study on Retriever Training
To demonstrate that using the well-trained RALM to guide the finetuning of the retriever is superior, we compared the retrieval results of the retriever finetuned with the raw LLM preferences to those finetuned with the preferences of RALM. We reported the recall rate of correct answers in the retrieved documents in Table VII.
The experimental results show that finetuning with the preferences of RALM generally outperforms finetuning with the raw LLM preferences. We observed that Replug - which evaluates document quality through answer probability - exhibited an unexpected recall rate decrease when using raw LLM, underperforming even the untuned baseline retriever. The instability arises because when the LLM can independently generate correct answers, using answer probability as a document quality metric becomes unreliable. Specifically, in such cases the LLM tends to assign inflated scores to all documents, making it impossible to discern their actual quality differences. In contrast, the well-trained RALM has a higher ability to discern documents and can more accurately differentiate between good and bad ones, thereby achieving better training outcomes. Meanwhile, the results also show that our model demonstrates leading performance across different recall metrics, with more pronounced improvements in R@5 and R@10, indicating that our model has greater advantages when retrieving multiple relevant documents.
TABLE VII: Experimental results on Recall rates(%) for retrievers trained with different methods. ”Baseline” refers to the retrieval results of the raw retriever and ”w/o RALM” indicates using the raw LLM without employing the well-trained RALM.
Method Avg.
[…]
BaselineBaseline 44.35 67.40 74.26
ReplugReplug 49.10 72.32 78.43
w/oRALMw/o\ RALM 43.82 67.25 74.25
LLMLLM-EmbedderEmbedder 49.45 71.48 77.78
w/oRALMw/o\ RALM 47.78 69.78 76.29
DACLDACL-RAGRAG 50.40 73.29 79.55
w/oRALMw/o\ RALM 48.55 70.71 77.33
TABLE VIII: Ablation studies for training stages in our DACL-RAG framework(%). We report the average EM and F1 scores here.
Training Stage Generator Training Retriever Training
Stage1 Stage2 Stage3 $uparrow EM\uparrow F1\uparrow$
$\checkmark$ 53.54 60.46 59.56 66.70
$\checkmark$ 55.66 63.15 59.83 67.01
$\checkmark$ 55.90 63.47 58.63 65.87
$\checkmark$ $\checkmark$ 57.56 64.72 60.04 67.17
$\checkmark$ $\checkmark$ 57.59 64.91 59.88 67.05
$\checkmark$ $\checkmark$ 57.44 64.52 60.14 67.31
$\checkmark$ $\checkmark$ $\checkmark$ 58.08 65.32 60.43 67.60
V-C3 Ablation Studies of Each Training Stage
Figure 8: Ablation studies for each training stage in our DACL-RAG framework. We demonstrate the EM score gains contributed by each training stage and compare with the results without curriculum learning strategies (w/o CL). ”With Retriever Training” employs the well-trained RALM (obtained during generator training) for both retriever training and final answer generation.
To systematically elucidate the contribution of each component to the overall performance, we conducted ablation studies in two aspects: ablation of individual training stages, and ablation of the curriculum learning strategy.
In the ablation study on training stages, we combined data from each stage and separately tested the performance of training the generator and the retriever. Specifically, we sampled from the training data of easy, common, and hard difficulty levels, keeping the total number of samples constant while varying the difficulty types contained in the samples, ensuring consistency between the samples and the number of backpropagation steps during training, and trained the model in order of difficulty. The results are shown in Table VIII.
In terms of generator training, we found that the absence of training data from any single stage led to suboptimal performance, and the model trained solely on easy data performed the worst. This indicates that in real retrieval scenarios, where retrieved documents vary in quality, using only easy data fails to train the model to resist interference from complex documents. Although training with common and hard difficulty data achieved relatively good performance, it was still not optimal compared to incorporating easy data, demonstrating the necessity of easy documents in training.
For retriever training, we observed that the model trained exclusively on hard samples performed the poorest. This is likely because the training samples contained only high-quality, less diverse documents, making it difficult for the model to develop strong discriminative ability. Similarly, after training with documents of varying difficulty levels, the model achieved optimal performance.
In parallel, we conducted an ablation study on the curriculum learning strategy. Specifically, we compared the model trained with the curriculum learning strategy (progressing stage by stage) against the model trained by directly feeding samples in random order. We demonstrated the improvement in model capability at each training stage, as well as the test results without using the curriculum learning strategy (w/o CL). The key findings are presented in Figure 8, where ”with RALM training” indicates training only the generator, and ”with retriever training” represents the results obtained by training the retriever on top of the pre-trained generator.
We observed that the performance of both the retriever and the generator steadily improves as the training stages progress. Furthermore, the removal of the curriculum learning strategy resulted in a significant performance decline. Notably, the model trained without curriculum learning for the retriever performed even worse than the model trained with only a single stage, indicating that neglecting the difficulty order of training samples leads to sub-optimal training outcomes.
V-C4 Case Study
@@
Table IX presents the case study comparing the document evaluation result of the raw LLM and the well-trained RALM. The result indicate that for some documents containing implicit knowledge required for answering questions, the raw LLM may generate erroneous judgments, whereas the well-trained RALM is more capable of distinguishing the quality of documents. This also highlights the necessity of the iterative process of training the RALM first and then training the retriever, as the well-trained RALM and the raw LLM may no longer share the same preferences for documents. Separate training may lead to suboptimal final outcomes.
TABLE IX: Case Study of document evaluation using raw LLM and well-trained RALM.
Question: Who scored a film based on a 1961 science fiction novel
by Stanislaw Lem?
@@
(Poland, Germany, and the Soviet Union). Franz Rottensteiner,
Lem’s former agent abroad, had this to say about Lem’s reception
on international markets: His best-known novels include ”Solaris”
(1961), ”His Master’s Voice” (”Głos pana”, 1968), and the late
”Fiasco” (”Fiasko”, 1987). ”Solaris” was made into a film in 1968
by Russian director Boris Nirenburg, a film in 1972 by Russian
director Andrei Tarkovsky—which won a Special Jury Prize at the
Cannes Film Festival in 1972—andan American film in 2002 by
Steven Soderbergh. ”Solaris” is not the only work of Lem’s
@@
VI Conclusion
In this study, we introduce DACL-RAG, a multi-level Data Augmentation strategy with a multi-stage Curriculum Learning paradigm for the RAG system. To the best of our knowledge, it’s one of the first time that the idea of human imitation learning is integrated with RAG training, which can effectively enhance the generalization and stability of the RAG system. Experiments on five open-domain question answering datasets on three different models provide substantial evidence of the framework’s effectiveness and generalization capability. Additionally, separate experiments conducted on the Retriever and the Generator have demonstrated the significant enhancements our method brings to each individual part.